RC Update and Run Plan for Aug 11, 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, remote

Tuesday Aug 11 2020.

General updates

  • Reasonable beam since yesterday. Beam taken away at 8am for RF recovery.
  • Dave Mack looking at BCM runs.
  • Target changed from 270->90 https://logbooks.jlab.org/entry/3824825
  • Hall A charge asymmetry feedback alarming overnight. It seems that outdated BCM gains were loaded. [1][2][3]
  • SoLID cerenkov moved from HMS to SHMS side. [4]
  • HMS certified to rotate down to 13.5 deg.
  • Angle encoder offsets from earthquake are fixed. [5]

Pending problem

Recent polarization measurements

  • 43% at 8am (270 deg) [6]
  • NMR at 90 deg [7]

Current Run Plan

Normal Production running

  • Prerequisites
    • Make sure the raster size is correct (4-4.5)
    • Monitor beam parameters/configuration very often and post an entry like this at the beginning of your shift and when run conditions have changed
    • Running on cell targets: raster always ON (4-4.5, size will be determined by the scraping checks), orbit locks ON, current ramp rate 1 muA/sec
    • Always take 120Hz of EDTM
    • Fill the BTA with realistic numbers. If the desired beam current is not delivered prorate your ABUs!
    • Change the IHWP every 4 hours when running on glass cells
  • Run plan
    • Re-establish beam in the Hall Short beam checkout
    • Production on 3He; if not able to reach requested current (27 uA), pro-rate the ABUs
    • Moller measurement during Wednesday Day shift (Murchana and Bill)

Notes for tomorrow's 7:45 meeting

Future

  • Opportunistic target polarization calibration and studies on Tuesday-Wednesday during beam off and Moller measurement (priority is beam)??
  • After ~ 10 days of running Hall C will go to 1 pass for ~3 days (plan for 17AUG)
  • EPR diode good for now, but check for future planned accesses.

Hall Access

When there is a down:

Notices in Effect for Shift Crew

  • Beam positions: [8]
  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If any of the Spectrometers magnets trip, take a screenshot of the PSU window and check the hall for smoke or water with the camera before resetting the interlock and post it to hclog.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[9]

Planned Accelerator Configuration Changes

  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links